Find the rectangular equation of the curve whose parametric equations are x = -2 cos(t) and y = 2 sin(t), 0° < t < 360°.
y² - x² = 1
x² + y² = 4
y² - x² = 4
x² + y² = 1

Answer by ikleyn(52781)

.
Find the rectangular equation of the curve whose parametric equations are x = -2 cos(t) and y = 2 sin(t), 0° < t < 360°.

y² - x² = 1
x² + y² = 4     <--- This and only this
y² - x² = 4
x² + y² = 1